OpenStudy (anonymous):

answer fully if u want medal!!! y is inversely proportional to x. Given that y=20 when x=2, find y when x=5.

OpenStudy (mathmale):

@sheelo: Beware: OpenStudy is not a source of free answers; you yourself must be involved in finding the solutions of the problems you post. What do you need to know to begin the solution of this problem?

OpenStudy (anonymous):

here x=k/y 1st find k if x=2 y=20 then so 2=k/20 =>k=40 now if x=5 and k=40 then 5=40/y =>y=40/5=8 so y=8

OpenStudy (mathmale):

Think: what does "inversely proportional" mean?

OpenStudy (mathmale):

@zaibali.qasmi : Thank you for trying to help. However, please read my previous statement. You could be of much more help to the person asking this question if you provide only enough info so that that person can solve this problem himself or herself. Do NOT do the work of others for them.

OpenStudy (anonymous):

rit...thanxx

OpenStudy (anonymous):

@mathmale ya i know how to make and equation but i could not understand how to find k

OpenStudy (anonymous):

now i got the idea how to find k thankyou both

OpenStudy (anonymous):

you welcome..

OpenStudy (anonymous):

@zaibali.qasmi r u pakistani?

OpenStudy (anonymous):

yes..n u..??

OpenStudy (mathmale):

I'm happy this conversation has worked out for you. Nevertheless, please do not ask for answers or offer medals for answers.

OpenStudy (anonymous):

me too from pak